# On the diversity of strongly-interacting Type IIn supernovae

### v1

@@ Acknowledgements.
The author acknowledges the help of Dr. T. Pitik in recreating Figure 22. This paper is supported by fundings from MIUR, PRIN 2017 (grant 20179ZF5KS). IS, AR, and NER are partially supported by the PRIN-INAF 2022 project “Shedding light on the nature of gap transients: from the observations to the models”. AR acknowledges the support GRAWITA Large Program Grant (PI P. D’Avanzo). YZC is supported by the National Natural Science Foundation of China (NSFC, Grant No. 12303054), the Yunnan Fundamental Research Projects (Grant No. 202401AU070063) and the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). JPA was funded by ANID, Millennium Science Initiative, ICN12_009. TP acknowledges the support by ANID through the Beca Doctorado Nacional 202221222222. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. LG, CPG and TEMB acknowledge financial support from the Spanish Ministerio de Ciencia e Innovación (MCIN), the Agencia Estatal de Investigación (AEI) 10.13039/501100011033, the European Social Fund (ESF) “Investing in your future”, the European Union Next Generation EU/PRTR funds, the Horizon 2020 Research and Innovation Programme of the European Union, and by the Secretary of Universities and Research (Government of Catalonia), under the PID2020-115253GA-I00 HOSTFLOWS project, 2021-SGR-01270, the 2019 Ramón y Cajal program RYC2019-027683-I, the 2021 Juan de la Cierva program FJC2021-047124-I, the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, and from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016, and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TWC acknowledges the Yushan Young Fellow Program by the Ministry of Education, Taiwan for the financial support. MB, EC and TP acknowledge the financial support from the Slovenian Research Agency (grants I0-0033, P1-0031, J1-8136, J1-2460 and Z1-1853) and the Young Researchers program. PC acknowledges support via Research Council of Finland (grant 340613). Based on observations collected at Copernico 1.82m telescope and Schmidt 67/92 telescope (Asiago Mount Ekar, Italy) INAF - Osservatorio Astronomico di Padova. Based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. Observations from the Nordic Optical Telescope were obtained through the NUTS2 collaboration, which are supported in part by the Instrument Centre for Danish Astrophysics (IDA). The data presented here were obtained in part with ALFOSC, which is provided by the Instituto de Astrofisica de Andalucia (IAA). This work makes use of data from the Las Cumbres Observatory network. The LCO team is supported by NSF grants AST–1911225 and AST– 1911151, and NASA SWIFT grant 80NSSC19K1639. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council. This work was based in part on observations made with the Italian Telescopio Nazionale Galileo (TNG), operated on the island of La Palma by the Fundación Galileo Galilei of the INAF (Istituto Nazionale di Astrofisica) at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias. The Gran Telescopio CANARIAS (GTC) is a 10.4m telescope with a segmented primary mirror. It is located in one of the top astronomical sites in the Northern Hemisphere: the Observatorio del Roque de los Muchachos (ORM, La Palma, Canary Islands). The GTC is a Spanish initiative led by the Instituto de Astrofísica de Canarias (IAC). The project is actively supported by the Spanish Government and the Local Government from the Canary Islands through the European Funds for Regional Development (FEDER) provided by the European Union. The project also includes the participation of Mexico (Instituto de Astronomía de la Universidad Nacional Autónoma de México (IA-UNAM) and Instituto Nacional de Astrofísica, Óptica y Electrónica (INAOE)), and the US University of Florida. Based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programmes 1103.D-0328 and 105.20TF.001. We acknowledge the use of public data from the Swift data archive. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST–1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. This work presents results from the European Space Agency (ESA) space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.References
@@
- Brennan et al. (2023) Brennan, S. J., Schulze, S., Lunnan, R., et al. 2023, arXiv e-prints, arXiv:2312.13280
@@ Appendix A Observations
Our target were monitored using the Schmidt and Copernico telescopes of the Asiago Observatory888https://www.oapd.inaf.it/sede-di-asiago/telescopes-and-instrumentations/, INAF (Italy); the Nordic Optical Telescope (NOT)999https://www.not.iac.es, the Telescopio Nazionale Galileo (TNG)101010https://www.tng.iac.es, the Liverpool Telescope (LT)111111https://telescope.livjm.ac.uk, and the Gran Telescopio CANARIAS (GTC)121212https://www.gtc.iac.es, all located in La Palma (Spain); the Rapid Eye Mount (REM)131313http://www.rem.inaf.it of INAF at La Silla and the Very Large Telescope (VLT)141414https://www.eso.org/public/teles-instr/paranal-observatory/vlt of ESO at Paranal (Chile). We also exploited the time allocated to other facilities to international collaborations, namely the NOT via NUTS2 (Nordic-optical-telescope Un-biased Transient Survey)151515https://nuts2.sn.ie/ and the ESO NTT161616https://www.eso.org/public/teles-instr/lasilla/ntt via ePESSTO+ (advanced Public ESO Spectroscopic Survey for Transient Objects, Smartt et al. 2015). These observations were complemented with data from transient surveys available from public archives, in particular, the Zwicky Transient Facility (ZTF, Bellm et al. 2019), the Las Cumbres Observatory (LCO, Brown et al. 2013), the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1, Chambers et al. 2019), the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018). Space observations obtained by the Gaia mission (Gaia Collaboration et al. 2018), and from NASA’s Neil Gehrels Swift Observatory (SWIFT, Roming et al. 2005) were also retrieved.
@@ B.1 Photometry
Photometric observations were reduced with standard techniques using IRAF171717https://iraf-community.github.io/ recipes and a number of different python packages, in particular astropy (Astropy Collaboration et al. 2022) and affiliated packages (astroquery, ccdproc, photutils). At first, the detector signature was removed with bias and flat-field corrections. Then, an algorithm for cosmic-ray rejection181818The algorithm is an implementation of the code described in van Dokkum (2001) as implemented by McCully et al. (2018). was applied. For astrometric and photometric calibration and for the measurement of the SN magnitudes the required recipes in the ecsnoopy package191919ecsnoopy is a python package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/ecsnoopy.html were implemented. In most cases (when the SN magnitude was below mag), the SN magnitude was measured after subtracting a template image from the observed frame, to better remove the contamination from the host galaxy. To this aim, reference archival images from public surveys such as PS1, SDSS (Sloan Digital Sky Survey, Kollmeier et al. 2019), or Skymapper (Keller et al. 2007) were used. With ecsnoopy, the registration of the template image was secured to the same pixel grid of the science image and then the code hotpants (Becker 2015) was used for the convolution of the two images to the same PSF and photometric scale. Finally, the instrumental magnitudes (or upper limits) were calibrated using photometric zero points measured from local stars with photometry retrieved from the public surveys mentioned above.
@@
Finally, for the near-infrared (NIR) observations, the same procedure as for the optical observations was used but including the preliminary subtraction of the sky background, which was obtained from the median combination of the dithered images for each filter. In this case, the nightly zero-points were computed with respect to the Two Micron All-Sky Survey (2MASS, Skrutskie et al. 2003) source catalog photometry. The magnitudes of the targets are reported in Tabs. 6-9, together with the magnitudes acquired from the public surveys mentioned above.
The available pre-supernova imaging was also checked for possible eruptive episodes in our SNe. In particular, through the available web tool202020https://fallingstar-data.com/forcedphot/, ATLAS forced photometry at the SN locations was obtained. In all cases, we concluded that there is no evidence of variation in the observed flux in the last decade. However, we stress that the limiting magnitude of ATLAS (20.0 mag in cyan and 19.5 mag in orange), corresponds to an absolute magnitude that ranges from to mag for the distance of our SNe, so the limit is not stringent and it is possible that any low-intensity precursor activity went undetected.
@@ B.2 Spectroscopy
The spectroscopic observations were in general reduced using standard prescriptions with the package foscgui212121foscgui is a python/pyraf-based graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. The spectra were corrected for bias, flat-field, and cosmics rejection, calibrated in wavelength in the 2D frame, and extracted to obtain the 1D spectrum. This was then calibrated in flux and corrected for second-order contamination (if required) and telluric absorptions, the latter with the aid of the spectrum of a hot spectrophotometric standard star. For the spectra taken using X-shooter, the publicly available data reduction pipeline EsoReflex222222https://www.eso.org/sci/software/esoreflex/ was used, following the same steps described above. The logs of spectroscopic observations are reported in Tabs. 10-13, together with the classification spectra retrieved from the Transient Name Server (TNS)232323https://www.wis-tns.org, when available.The spectral evolution of the targets through the most significant spectra is shown in Figs. 4, 5, 7, and 8.

### v2

@@ 5 Conclusions
Data availability
Tables with photometric measurements, spectral log of observations, and fit on the H and H are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We thank the anonymous referee for their helpful insights. The author acknowledges the help of Dr. T. Pitik in recreating Figure 22. This paper is supported by fundings from MIUR, PRIN 2017 (grant 20179ZF5KS). IS, AR, and NER are partially supported by the PRIN-INAF 2022 project “Shedding light on the nature of gap transients: from the observations to the models”. AR acknowledges the support GRAWITA Large Program Grant (PI P. D’Avanzo). YZC is supported by the National Natural Science Foundation of China (NSFC, Grant No. 12303054), the National Key Research and Development Program of China (Grant No. 2024YFA1611603), the Yunnan Fundamental Research Projects (Grant No. 202401AU070063), and the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). JPA was funded by ANID, Millennium Science Initiative, ICN12_009. TP acknowledges the support by ANID through the Beca Doctorado Nacional 202221222222. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. LG, CPG and TEMB acknowledge financial support from the Spanish Ministerio de Ciencia e Innovación (MCIN), the Agencia Estatal de Investigación (AEI) 10.13039/501100011033, the European Social Fund (ESF) “Investing in your future”, the European Union Next Generation EU/PRTR funds, the Horizon 2020 Research and Innovation Programme of the European Union, and by the Secretary of Universities and Research (Government of Catalonia), under the PID2020-115253GA-I00 HOSTFLOWS project, 2021-SGR-01270, the 2019 Ramón y Cajal program RYC2019-027683-I, the 2021 Juan de la Cierva program FJC2021-047124-I, the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme, and from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016, and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TWC acknowledges the Yushan Young Fellow Program by the Ministry of Education, Taiwan for the financial support. MB, EC and TP acknowledge the financial support from the Slovenian Research Agency (grants I0-0033, P1-0031, J1-8136, J1-2460 and Z1-1853) and the Young Researchers program. PC acknowledges support via Research Council of Finland (grant 340613). MS is funded by the Independent Research Fund Denmark (IRFD, grant number 10.46540/2032-00022B). Based on observations collected at Copernico 1.82m telescope and Schmidt 67/92 telescope (Asiago Mount Ekar, Italy) INAF - Osservatorio Astronomico di Padova. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101004719. This material reflects only the authors views and the Commission is not liable for any use that may be made of the information contained therein. Based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. Observations from the Nordic Optical Telescope were obtained through the NUTS2 collaboration, which are supported in part by the Instrument Centre for Danish Astrophysics (IDA). The data presented here were obtained in part with ALFOSC, which is provided by the Instituto de Astrofisica de Andalucia (IAA). This work makes use of data from the Las Cumbres Observatory network. The LCO team is supported by NSF grants AST–1911225 and AST– 1911151, and NASA SWIFT grant 80NSSC19K1639. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council. This work was based in part on observations made with the Italian Telescopio Nazionale Galileo (TNG), operated on the island of La Palma by the Fundación Galileo Galilei of the INAF (Istituto Nazionale di Astrofisica) at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias. The Gran Telescopio CANARIAS (GTC) is a 10.4m telescope with a segmented primary mirror. It is located in one of the top astronomical sites in the Northern Hemisphere: the Observatorio del Roque de los Muchachos (ORM, La Palma, Canary Islands). The GTC is a Spanish initiative led by the Instituto de Astrofísica de Canarias (IAC). The project is actively supported by the Spanish Government and the Local Government from the Canary Islands through the European Funds for Regional Development (FEDER) provided by the European Union. The project also includes the participation of Mexico (Instituto de Astronomía de la Universidad Nacional Autónoma de México (IA-UNAM) and Instituto Nacional de Astrofísica, Óptica y Electrónica (INAOE)), and the US University of Florida. Based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programmes 1103.D-0328 and 105.20TF.001. We acknowledge the use of public data from the Swift data archive. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST–1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. This work presents results from the European Space Agency (ESA) space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.References
@@
- Brennan et al. (2024) Brennan, S. J., Schulze, S., Lunnan, R., et al. 2024, A&A, 690, A259
@@
- Habergham et al. (2014) Habergham, S. M., Anderson, J. P., James, P. A., & Lyman, J. D. 2014, MNRAS, 441, 2230
- Hiramatsu et al. (2024) Hiramatsu, D., Berger, E., Gomez, S., et al. 2024, arXiv e-prints, arXiv:2411.07287
@@
- Ransome et al. (2022) Ransome, C. L., Habergham-Mawson, S. M., Darnley, M. J., James, P. A., & Percival, S. M. 2022, MNRAS, 513, 3564
- Ransome & Villar (2024) Ransome, C. L. & Villar, V. A. 2024, arXiv e-prints, arXiv:2409.10596
@@
- Rodríguez et al. (2023) Rodríguez, Ó., Maoz, D., & Nakar, E. 2023, ApJ, 955, 71
@@ Appendix A Observations
Our targets were monitored using the Schmidt and Copernico telescopes of the Asiago Observatory111111https://www.oapd.inaf.it/sede-di-asiago/telescopes-and-instrumentations/, INAF (Italy); the Nordic Optical Telescope (NOT)121212https://www.not.iac.es, the Telescopio Nazionale Galileo (TNG)131313https://www.tng.iac.es, the Liverpool Telescope (LT)141414https://telescope.livjm.ac.uk, and the Gran Telescopio CANARIAS (GTC)151515https://www.gtc.iac.es, all located in La Palma (Spain); the Rapid Eye Mount (REM)161616http://www.rem.inaf.it of INAF at La Silla and the Very Large Telescope (VLT)171717https://www.eso.org/public/teles-instr/paranal-observatory/vlt of ESO at Paranal (Chile). We also exploited the time allocated to other facilities to international collaborations, namely the NOT via NUTS2 (Nordic-optical-telescope Un-biased Transient Survey)181818https://nuts2.sn.ie/ and the ESO NTT191919https://www.eso.org/public/teles-instr/lasilla/ntt via ePESSTO+ (advanced Public ESO Spectroscopic Survey for Transient Objects, Smartt et al. 2015). These observations were complemented with data from transient surveys available from public archives, in particular, the Zwicky Transient Facility (ZTF, Bellm et al. 2019), the Las Cumbres Observatory (LCO, Brown et al. 2013), the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1, Chambers et al. 2019), the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018). Space observations obtained by the Gaia mission (Gaia Collaboration et al. 2018), and from NASA’s Neil Gehrels Swift Observatory (SWIFT, Roming et al. 2005) were also retrieved.
@@ B.1 Photometry
Photometric observations were reduced with standard techniques using IRAF202020https://iraf-community.github.io/ recipes and a number of different python packages, in particular astropy (Astropy Collaboration et al. 2022) and affiliated packages (astroquery, ccdproc, photutils). At first, the detector signature was removed with bias and flat-field corrections. Then, an algorithm for cosmic-ray rejection212121The algorithm is an implementation of the code described in van Dokkum (2001) as implemented by McCully et al. (2018). was applied. For astrometric and photometric calibration and for the measurement of the SN magnitudes the required recipes in the ecsnoopy package222222ecsnoopy is a python package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/ecsnoopy.html were implemented. In most cases (when the SN magnitude was below mag), the SN magnitude was measured after subtracting a template image from the observed frame, to better remove the contamination from the host galaxy. To this aim, reference archival images from public surveys such as PS1, SDSS (Sloan Digital Sky Survey, Kollmeier et al. 2019), or Skymapper (Keller et al. 2007) were used. With ecsnoopy, the registration of the template image was secured to the same pixel grid of the science image and then the code hotpants (Becker 2015) was used for the convolution of the two images to the same PSF and photometric scale. Finally, the instrumental magnitudes (or upper limits) were calibrated using photometric zero points measured from local stars with photometry retrieved from the public surveys mentioned above.
@@
Finally, for the near-infrared (NIR) observations, the same procedure as for the optical observations was used but including the preliminary subtraction of the sky background, which was obtained from the median combination of the dithered images for each filter. In this case, the nightly zero-points were computed with respect to the Two Micron All-Sky Survey (2MASS, Skrutskie et al. 2003) source catalogue photometry. The magnitudes of the targets are available online only (see the Data availability section), together with the magnitudes acquired from the public surveys mentioned above.
The available pre-supernova imaging was also checked for possible eruptive episodes in our SNe. In particular, through the available web tool232323https://fallingstar-data.com/forcedphot/, ATLAS forced photometry at the SN locations was obtained. In all cases, we concluded that there is no evidence of variation in the observed flux in the last decade. However, we stress that the limiting magnitude of ATLAS (20.0 mag in cyan and 19.5 mag in orange), corresponds to an absolute magnitude that ranges from to mag for the distance of our SNe, so the limit is not stringent and it is possible that any low-intensity precursor activity went undetected.
@@ B.2 Spectroscopy
The spectroscopic observations were in general reduced using standard prescriptions with the package foscgui242424foscgui is a python/pyraf-based graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. The spectra were corrected for bias, flat-field, and cosmics rejection, calibrated in wavelength in the 2D frame, and extracted to obtain the 1D spectrum. This was then calibrated in flux and corrected for second-order contamination (if required) and telluric absorptions, the latter with the aid of the spectrum of a hot spectrophotometric standard star. For the spectra taken using X-shooter, the publicly available data reduction pipeline EsoReflex252525https://www.eso.org/sci/software/esoreflex/ was used, following the same steps described above. The logs of spectroscopic observations are available online only (see the Data availability section). The spectral evolution of the targets through the most significant spectra is shown in Figs. 4, 5, 7, and 8.